Lactic acid solution is formed

-write reaction where lactic acid is reacted with water

-given Ka=1.4 x 10-^4 and [lactic acid] = 0.20 M determine [H3O+]

-determine Kb for lactic acid

-determine [OH-]


1
Expert's answer
2021-08-20T01:52:49-0400

conjugate acid is lactic acid

Ka of lactic acid = 1.4 x 10-4

Kb = Kw / Ka = 10-14 / (1.4 x 10-4) = 7.14 x 10-11

Kb = [OH-] [lactic acid] / [lactate]

[OH-] = [(7.14 x 10-11) (0.25)]1/2 = 4.2 x 10-6